\begin{thebibliography}{99}
%%
%\bibitem{AJY}
%	Achleitner, F., J\"ungel, A., Yamamoto, M.,
%	Large-time asymptotics of a fractional drift-diffusion-Poisson system via the entropy method,
%	preprint.
%%
%\bibitem{Bl-Dl}
%	Biler, P., Dolbeault, J.,
%	Long time behavior of solutions to Nernst-Planck and Debye-H\"{u}ckel 
%drift-diffusion systems,
%	Ann. Henri Poincar\'{e} {\bf 1} (2000), 461--472.\\
%	{\tt doi:10.1007/s000230050003}
%%
%
\bibitem{Blmntr-Gtr}
	Blumenthal, R.M., Getoor, R.K.,
	Some theorems on stable processes,
	Trans. Amer. Math. Soc., {\bf 95} (1960), 263--273.
%
%
\bibitem{Brndls}
	Brandolese, L.,
	Space-time decay of Navier-Stokes flows invariant under rotations,
	Math. Ann. {\bf 329} (2004), 685--706
%
\bibitem{Brndls-Vgnrn}
	Brandolese, L., Vigneron, F.,
	New asymptotic profiles of nonstationary solutions of the Navier-Stokes system,
	J. Math. Pures Appl. {\bf 88} (2007), 64--86.
%
\bibitem{Brndls-Krch}
	Brandolese, L., Karch, G.,
	Far field asymptotics of solutions to convection equation with anomalous diffusion,
	J. Evol. Equ. {\bf 8} (2008), 307--326.
%%
%\bibitem{Br-Cl}
%	Bournaveas, N., Calvez, V.,
%	The one-dimensional Keller-Segel model with fractional diffusion of cells,
%	Nonlinearity {\bf 23} (2010), 923--935.\\
%	{\tt doi:10.1088/0951-7715/23/4/009}
%%
\bibitem{Cffrll-Vssur}
	Caffarelli, L.A., Vasseur, A.,
	Drift diffusion equations with fractional diffusion and the quasi-geostrophic equation,
	Ann. of Math. (2) {\bf 171} (2010), 1903--1930.
%	{\tt doi:10.4007/annals.2010.171.1903}
%%
%\bibitem{Cr}
%	Carpio, A.,
%	Large-time behavior in incompressible Navier-Stokes equation,
%	SIAM J. Math. Anal. {\bf 27} (1996), 449--475.
%%	{\tt doi:10.1137/S0036141093256782}
%%
\bibitem{Ch-L}
	Chae, D., Lee, J.,
	Global well-posedness in the super-critical dissipative quasi-geostrophic equations,
	Comm. Math. Phys. {\bf 233} (2003), 297--311.
%	{\tt doi:10.1007/s00220-002-0750-z}
%
\bibitem{C-C-W}
	Constantin, P., Cordoba, D., Wu, J.,
	On the critical dissipative quasi-geostrophic equation,
	Indiana Univ. Math. J. {\bf 50} (2001), 97--107.
%	{\tt doi:10.1512/iumj.2001.50.2153}
%
\bibitem{Cnstntn-Mjd-Tbk}
	Constantin, P., Majda, A., Tabak, E.,
	Formation of strong fronts in the 2-D quasi-geostrophic thermal active scalar,
	Nonlinearity {\bf 7} (1994), 1495-1533.
%	{\tt doi:10.1088/0951-7715/7/6/001}
%%
\bibitem{Crdb-Crdb}
	Cordoba, A., Cordoba, D.,
	A maximum principle applied to quasi-geostrophic equations,
	Comm. Math. Phys. {\bf 249} (2004), 511--528.
%	{\tt doi:10.1007/s00220-004-1055-1}
%%
%\bibitem{D-D-L}
%	Dong, H., Du, D., Li, D.,
%	Finite time singularities and global well-posedness for fractal Burgers equations,
%	Indiana Univ. Math. J. {\bf 58} (2009), 807--821.\\
%	{\tt doi:10.1512/iumj.2009.58.3505}
%
\bibitem{E-Z}
	Escobedo, M., Zuazua, E.,
	Large time behavior for convection-diffusion equation in $\mathbb{R}^n$,
	J. Funct. Anal. {\bf 100} (1991), 119--161.
%	{\tt doi:10.1016/0022-1236(91)90105-E}
%
%%
%\bibitem{F-M}
%	Fujigaki, Y., Miyakawa, T.,
%	Asymptotic profiles of nonstationary incompressible Navier-Stokes flows 
%in the whole space,
%	SIAM J. Math. Anal. {\bf 33} (2001), 523--544.\\
%	{\tt doi:10.1137/S0036141000367072}
%%
\bibitem{H-Y-Z}
	Hajaiej, H., Yu, X., Zhai, Z.,
	Fractional Gagliardo-Nirenberg and Hardy inequalities under Lorentz norms,
	J. Math. Anal. Appl. {\bf 396} (2012), 569--577.
%	{\tt doi:10.1016/j.jmaa.2012.06.054}
%%
%\bibitem{H-K-N}
%	Hayashi, N., Kaikina, E.I., Naumkin, P.I.,
%	Large time asymptotics for the BBM-Burgers equation,
%	Ann. Henri Poincar\'e {\bf 8} (2007), 485--511.\\
%	{\tt doi:10.1007/s00023-006-0314-4}
%%
%\bibitem{H-N}
%	Hayashi, N., Naumkin, P.I.,
%	Asymptotics for the Korteweg-de Vries-Burgers equation,
%	Acta Math. Sin. {\bf 22} (2006), 1441--1456.\\
%	{\tt doi:10.1007/s10114-005-0677-3}
%%
\bibitem{Hrmndr}
	H\"ormander, L.,
	Estimates for translation invariant operators in $L^p$ spaces,
	Acta Math. {\bf 104} (1960), 93--140.
%%
%\bibitem{I-I-K}
%	Ishige, K., Ishiwata, M., Kawakami, T.,
%	The decay of the solutions for the heat equation with a potential,
%	Indiana Univ. Math. J. {\bf 58} (2009), 2673--2708.\\
%	{\tt doi:10.1512/iumj.2009.58.3771}
%%
%\bibitem{I-K}
%	Ishige, K., Kawakami, T.,
%	Refined asymptotic profiles for a semilinear heat equation,
%	Math. Ann. {\bf 353} (2012), 161--192.\\
%	{\tt doi:10.1007/s00208-011-0677-9}
%%
%\bibitem{I-K-K}
%	Ishige, K., Kawakami, T., Kobayashi, K.,
%	Asymptotics for a nonlinear integral equation with a generalized heat kernel,
%	J. Evol. Equ. {\bf 14} (2014), 749--777.\\
%	{\tt doi:10.1007/s00028-014-0237-3}
%%%
%\bibitem{I-K-M}
%	Ishige, K., Kawakami, M., Michihisa, H.,
%	Asymptotic expansions of solutions of fractional diffusion equations,
%	SIAM J. Math. Anal. {\bf 49} (2017), 2167--2190.
%%
%\bibitem{I}
%	Iwabuchi, T.,
%	Global solutions for the critical Burgers equation in the Besov spaces and the large time behavior,
%	Ann. Inst. H. Poincar\'e Anal. Non Lin\'eaire {\bf 32} (2015), 687--713.\\
%	{\tt doi:10.1016/j.anihpc.2014.03.002}
%%
\bibitem{J05}
	Ju, N.,
	The maximum principle and the global attractor for the dissipative 2D quasi-geostrophic equations,
	Comm. Math. Phys. {\bf 255} (2005), 161--181.
%	{\tt doi:10.1007/s00220-004-1256-7}
%
\bibitem{J04}
	Ju, N.,
	Existence and uniqueness of the solution to the dissipative 2D quasi-geostrophic equations in the Sobolev space,
	Comm. Math. Phys. {\bf 251} (2004), 365--376.
%	{\tt doi:10.1007/s00220-004-1062-2}
%%
%\bibitem{Jngl}
%	J\"ungel, A.,
%	Qualitative behavior of solutions of a degenerate nonlinear drift-diffusion model for semiconductors,
%	Math. Models Methods Appl. Aci. {\bf 5} (1995), 497--518.\\
%	{\tt doi:10.1142/S0218202595000292}
%%
%\bibitem{K-RP}
%	Kaikina, E.I., Ruiz-Paredes, H.F.,
%	Second term of asymptotics for KdVB equation with large initial data,
%	Osaka J. Math. {\bf 42} (2005), 407--420.
%%
\bibitem{K-PP}
	Kalamajska, A., Pietruska-Paluba, K.,
	Gagliardo Nirenberg inequalities in logarithmic spaces,
	Colloq. Math. {\bf 106} (2006), 93--107.
%	{\tt doi:10.4064/cm106-1-8}
%
%%
%\bibitem{KtM_B}
%	Kato, M.,
%	Large time behavior of solutions to the generalized Burgers equations,
%	Osaka J. Math. {\bf 44} (2007), 923--943.
%%
%\bibitem{KtM}
%	Kato, M.,
%	Sharp asymptotics for a parabolic system of chemotaxis in one space dimension,
%	Differential Integral Equations {\bf 22} (2009), 35--51.
%%
\bibitem{K-P}
	Kato, T., Ponce, G.,
	Commutator estimates and the Euler and Navier-Stokes equations,
	Comm. Pure Appl. Math. {\bf 41} (1988), 891--907.
%	{\tt doi:10.1002/cpa.3160410704}
%
\bibitem{K-N-V}
	Kiselev, A., Nazarov, F., Volberg, A.,
	Global well-posedness for the critical 2D dissipative quasi-geostrophic equation,
	Invent. Math. {\bf 167} (2007), 445-453.
%
\bibitem{K-K}
	Kobayashi, R., Kawashima, S., 
	Decay estimates and large time behavior of solutions to the 
    drift-diffusion system,
	Funkcial. Ekvac. {\bf 51} (2008), 371--394.
%	{\tt doi:10.1619/fesi.51.371}
%
\bibitem{Klkltsv}
	Kolokoltsov, V.N.,
	Markov processes, semigroups and generators,
	De Gruyter, Berlin, 2011.
%
%%
%\bibitem{K-O}
%	Kurokiba, M., Ogawa, T.,
%	Wellposedness for the drift-diffusion system in
%	$L^p$
%	arising from the semiconductor device simulation,
%	J. Math. Anal. Appl. {\bf 342} (2008), 1052--1067.\\
%	{\tt doi:10.1016/j.jmaa.2007.11.017}
%%%
%\bibitem{L-Rdrg}
%	Li, D., Rodorigo, J.,
%	Blow up for the generalized surface quasi-geostrophic equation with supercritical dissipation,
%	Comm. Math. Phys. {\bf 286} (2009), 111-124.
%%\bibitem{L-L-Z}
%	Li, D., Rodrigo, J.L., Zhang, X.,
%	Exploding solutions for a nonlocal quadratic evolution problem,
%	Rev. Mat. Iberoamericana {\bf 26} (2010), 295--332.\\
%	{\tt doi:10.4171/RMI/602}
%
\bibitem{L-S}
	Liskevich, V.A., Semenov, Y.A.,
	Some problems on Markov semigroups,
	Schr\"odinger operators, Markov semigroups, wavelet analysis, operator algegras, 163--217,
	Math. Top., {\bf 11}, Akademie Verlag, Berlin, 1996.
%
%\bibitem{M-T}
%	Matsumoto, T., Tanaka, N.,
%	Lipschitz semigroup approach to drift-diffusion systems,
%	RIMS K\^oky\^uroku Bessatsu {\bf B15} (2009), 147--177.
%%
%\bibitem{M-W}
%	Miao, C., Wu, G.,
%	Global well-posedness of the critical Burgers equation in critical Besov spaces,
%	J. Differential Equations {\bf 247} (2009), 1673--1693.\\
%	{\tt doi:10.1016/j.jde.2009.03.028}
%%
\bibitem{Mkhln}
	Mikhlin, S.G.,
	Multidimensional singular integrals and integral equations,
	Pergamon Press, Oxford, New York, Paris, 1965.
%
\bibitem{Mur06}
	Miura, H.,
	Dissipative quasi-geostrophic equation for large initial data in the critical Sobolev space,
	Comm. Math. Phys. {\bf 267} (2006), 141--157.
%
%
\bibitem{Mykw00}
	Miyakawa, T.,
	On space-time decay properties of nonstationary incompressible Navier-Stokes flows in $R^n$,
	Funkcial. Ekvac. {\bf 43} (2000), 541--557.
%
\bibitem{Mykw-Schnbk}
	Miyakawa, T., Schonbek, M.,
	On optimal decay rates for weak solutions to the Navier-Stokes equations in $\mathbb{R}^n$,
	Math. Bohemica, {\bf 126} (2001), 443--455.
%
\bibitem{M-N-S-S}
	Mizuta, Y., Nagai, E., Sawano, Y., Shimomura, T.,
	Gagliardo-Nirenberg inequality for generalized Riesz potentials of functions in Musielak-Orlicz spaces,
	Arch. Math. {\bf 98} (2012), 253--263.
%	{\tt doi:10.1007/s00013-012-0362-6}
%
%%
%\bibitem{M}
%	Mock, M. S.,
%	Asymptotic behavior of solutions
%	of  transport equations for semiconductor devices,
%	J. Math. Anal. Appl. {\bf 49} (1975), 215--225.\\
%	{\tt doi:10.1016/0022-247X(75)90172-9}
%%
%%
%\bibitem{N-S-U}
%	Nagai, T., Syukuinn, R., Umesako, M.,
%	Decay property and asymptotic profiles of bounded solutions
%	to a parabolic system of chemotaxis in $\mathit R^n$,
%	Funkcial. Ekvac. {\bf 46} (2003), 383--407.\\
%	{\tt doi:10.1619/fesi.46.383}
%
%\bibitem{N-Y}
%	Nagai, T., Yamada, T.,
%	Large time behavior of bounded solutions to a parabolic system of chemotaxis in the whole space,
%	J. Math. Anal. Appl. {\bf 336} (2007), 704--726.\\
%	{\tt doi:10.1016/j.jmaa.2007.03.014}
%%
%\bibitem{N-S_Mono}
%	Naumkin, P.I., Shishmar\"ev, I.A.,
%	Nonlinear nonlocal equations in the theory of waves,
%	Translations of Mathematical Monograph {\bf 133},
%	American Mathematical Society, Providence, RI, 1994.
%%
%\bibitem{N-S_DE}
%	Naumkin, P.I., Shishmar\"ev, I.A.,
%	An asymptotic relationship between solutions of different nonlinear equations for large time values. I, II,
%	Differential Equations {\bf 30} (1994), 806--814, 1329--1340.
%%
\bibitem{Og-Y}
	Ogawa, T., Yamamoto, M.,
	Asymptotic behavior of solutions to drift-diffusion system with generalized dissipation,
	Math. Models Methods Appl. Sci. {\bf 19} (2009), 939-967.
%
%\bibitem{O-S}
%	Ogawa, T., Shimizu, S.,
%	The drift-diffusion system in two-dimensional critical Hardy space,
%	J. Funct. Anal. {\bf 255} (2008), 1107--1138.\\
%	{\tt doi:10.1016/j.jfa.2008.05.020}
%
\bibitem{S-S}
	Shibata, Y., Shimizu, S.,
	A decay property of the Fourier transform and its application to the Stokes problem,
	J. Math. Fluid Mech. {\bf 3} (2001), 213--230.\\
	{\tt doi:10.1007/PL00000970}
%%%
\bibitem{S}
	Stein, E.M.,
	Singular Integrals and Differentiability Properties of Functions,
	Princeton University Press,
	Princeton, New Jersey, 1970.
%%
%\bibitem{S-Y-K}
%	Sugiyama, Y., Yamamoto, M., Kato, K.,
%	Local and global solvability and blow up for the drift-diffusion equation with the fractional dissipation with the critical space,
%	J. Differential Equations {\bf 258} (2015), 2983--3010.\\
%	{\tt doi:10.1016/j.jde.2014.12.033}
%%%
%\bibitem{Yd}
%	Yamada, T.,
%	Higher-order asymptotic expansions for a parabolic system modeling chemotaxis in the whole space,
%	Hiroshima Math. J. {\bf 39} (2009), 363--420.
%%
%\bibitem{Y}
%	Yamamoto, M.,
%	Large time behavior of solutions to the drift-diffusion equation with fractional dissipation,
%	Differential Integral Equations {\bf 25} (2012), 731--758.
%%
%\bibitem{Ym}
%	Yamamoto, M.,
%	Asymptotic expansion of solutions to the drift-diffusion equation with large initial data,
%	J. Math. Anal. Appl. {\bf 369} (2010), 144--163.\\
%	{\tt doi:10.1016/j.jde.2014.12.033}
%%
%\bibitem{Y-S-K}
%	Yamamoto, M., Kato, K., Sugiyama, Y.,
%	Existence and analyticity of solutions to the drift-diffusion equation with critical dissipation,
%	Hiroshima Math. J. {\bf 44} (2014), 275-313.
%%
%\bibitem{Y-S}
%	Yamamoto, M., Sugiyama, Y.,
%	Asymptotic behavior of solutions to the drift-diffusion equation with critical dissipation,
%	Annales Henri Poincar\'e {\bf 17} (2016), 1331--1352.\\
%	{\tt doi:10.1007/s00023-015-0428-7}
%%
\bibitem{Y-S-NA}
	Yamamoto, M., Sugiyama, Y.,
	Asymptotic expansion of solutions to the drift-diffusion equation with fractional dissipation,
	Nonlinear Anal. {\bf 141} (2016), 57-87.
%
\bibitem{Y-S18aX}
	Yamamoto, M., Sugiyama, Y.,
	Spatial-decay of solutions to the quasi-geostrophic equation with the critical and the supercritical dissipation,
	to appear in Nonlinearity,
	{\tt arXiv:1807.06793v1}.
%
 \bibitem{Z}
 	Ziemer, W.P.,
 	Weakly Differentiable Functions,	
 	Graduate Texts in Math., {\bf 120},
 	Springer Verlag,
 	New York,
 	1989.
\end{thebibliography}

\documentclass[11pt,draft,dvips]{amsart}  
\usepackage{amssymb,amsmath,amsfonts}  
\usepackage[dvips]{graphicx,color}
\usepackage{cite}
\allowdisplaybreaks[1]
\definecolor{pgray}{gray}{0.8}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{cor}[theorem]{Corollary}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{lemma}[theorem]{Lemma}

\numberwithin{equation}{section}
\title{\mbox{}}

\begin{document}
\begin{center}
{\bf \Large{
	Optimal Estimates for Far Field Asymptotics of Solutions to the Quasi-Geostrophic Equation
}}\\
\vspace{5mm}
{\sc\large
	Masakazu Yamamoto}\footnote{Graduate School of Science and Technology,
	Niigata University,
	Niigata 950-2181, Japan}\qquad
{\sc\large
	Yuusuke Sugiyama}\footnote{Department of Engineering,
	The University of Shiga Prefecture,
	Hikone 522-8533, Japan}
\end{center}
\maketitle
\vspace{-15mm}
\begin{abstract}
The initial value problem for the two dimensional dissipative quasi-geostrophic equation of the critical and the supercritical cases is considered.
Anomalous diffusion on this equation provides slow decay of solutions as the spatial parameter tends to infinity.
In this paper, uniform estimates for far field asymptotics of solutions are given.
\end{abstract}
%

\section{Introduction}
We derive far field asymptotics of solutions of the following initial-value problem:
\begin{equation}\label{qg}
\left\{
\begin{array}{lr}
	\partial_t \theta + (-\Delta)^{\alpha/2} \theta + \nabla \cdot (\theta\nabla^\bot\psi) = 0,
	&
	t > 0,~ x \in \mathbb{R}^2,\\
	(-\Delta)^{1/2}\psi = \theta,
	&
	t > 0,~ x \in \mathbb{R}^2,\\
	\theta (0,x) = \theta_0 (x),
	&
	x \in \mathbb{R}^2,
\end{array}
\right.
\end{equation}
where $\partial_t = \partial/\partial t,~ \nabla^\bot = (-\partial_2, \partial_1),~ \partial_j = \partial/\partial x_j$ for $j = 1,2$, and $(-\Delta)^{\alpha/2} \varphi = \mathcal{F}^{-1} [|\xi|^\alpha \mathcal{F} [\varphi]]$ for $0 < \alpha \le 2$.
The unknown function $\theta = \theta (t,x)$ stands for the potential temperature and $\psi = \psi (t,x)$ is the stream function (cf.\cite{Cnstntn-Mjd-Tbk}).
The fluid velocity is represented by $\nabla^\bot \psi = (-R_2 \theta, R_1 \theta)$ and $R_j = \partial_j (-\Delta)^{-1/2}$ is the Riesz transform.
When $\alpha = 1$ and $0 < \alpha < 1$, the scaling property of \eqref{qg} is in the critical and the supercritical, respectively.
In those cases, it is well-known that some smallness and smoothness for the initial-data are required to obtain the global existence of solutions in time.
Global existence of solutions in scale-invariant spaces is important problem also in the study for the Navier-Stokes flow.
Especially the critical quasi-geostrophic equation has similar stracture as the Navier-Stokes equations.
Furthermore, in the critical and the supercritical cases, the quasi-geostrophic equation seems to be elliptic and hyperbolic, respectively.
Hence the several methods for parabolic equations are not working for \eqref{qg}.
Because of those reasons, the quasi-geostrophic equation is considered by many authors (see \cite{Cffrll-Vssur,Ch-L,C-C-W,Crdb-Crdb,J05,J04,K-N-V,Mur06}).
In this paper, we treat the global solution in time which satisfies that
\begin{equation}\label{exist}
	\theta \in C \bigl( [0,\infty), L^1 (\mathbb{R}^2) \cap L^\infty (\mathbb{R}^2) \bigr).
\end{equation}
Then the mass-conservation and the uniform decay in time hold:
\begin{equation}\label{mass}
	\int_{\mathbb{R}^2} \theta (t,x) dx
	=
	\int_{\mathbb{R}^2} \theta_0 (x) dx,
	\quad
	\left\| \theta (t) \right\|_{L^\infty (\mathbb{R}^2)}
	\le
	C (1+t)^{-2/\alpha}.
\end{equation}
Those properties are confirmed initial data that is small and smooth.
In the recent paper \cite{Y-S18aX}, the smoothness and the upper bound of spatial decay of the solution are proved.
Namely, upon the condition $\theta_0 \in H^\sigma (\mathbb{R}^2)$ for $\sigma > 2$ and $|x|^2 \theta_0 \in L^q (\mathbb{R}^2)$ for $q > 2/\alpha$, the solution satisfies that
\begin{equation}\label{drbdd}
	\bigl\| (-\Delta)^{\sigma/2} \theta (t) \bigr\|_{L^2 (\mathbb{R}^2)}
	\le
	C (1+t)^{-\frac1\alpha-\frac\sigma\alpha}
\end{equation}
and
\begin{equation}\label{wttheta}
	\bigl\| |x|^2 \theta (t) \bigr\|_{L^q (\mathbb{R}^2)}
	\le
	C (1+t)^{\frac{2}{\alpha q}}.
\end{equation}
Those estimates are optimal since the fundamental solution $G_\alpha (t,x) = (2\pi)^{-1} \mathcal{F}^{-1} [e^{-t |\xi|^\alpha}] (x)$ of the linear equation $\partial_t \theta + (-\Delta)^{\alpha/2} \theta = 0$ fulfills that $\| (-\Delta)^{\sigma/2} G_\alpha (t) \|_{L^2 (\mathbb{R}^2)} = t^{-\frac1\alpha-\frac\sigma\alpha} \| (-\Delta)^{\sigma/2} G_\alpha (1) \|_{L^2 (\mathbb{R}^2)}$ and $\| |x|^2 G_\alpha (t) \|_{L^q (\mathbb{R}^2)}$ $= t^{\frac{2}{\alpha q}} \| |x|^2 G_\alpha (1) \|_{L^q (\mathbb{R}^2)}$.
When $\alpha = 1$, this fundamental solution is given by the Poisson kernel $P$.
Moreover the lower bound of spatial decay is derived in \cite{Y-S18aX}:
\begin{equation}\label{L2}
	\bigl\| |x|^2 \bigl( \theta (t) - M G_\alpha (t) \bigr) \bigr\|_{L^2 (\mathbb{R}^2)}
	\le
	C L_\alpha (t) \bigl( \log (2+t) \bigr)^{1/2},
\end{equation}
where $M = \int_{\mathbb{R}^2} \theta_0 (x) dx$, and $L_\alpha (t) = \log (2+t)$ for $\alpha = 1$ and $L_\alpha (t) = 1$ for $0 < \alpha < 1$.
Since $\| |x|^2 G_\alpha (t) \|_{L^2 (\mathbb{R}^2)} = + \infty$ and $G_\alpha \in C ((0,\infty) \times \mathbb{R}^2)$, \eqref{L2} intends that $\theta$ and $M G_\alpha$ are canceled in far field.
Therefore the asymptotic profile of $\theta$ as $|x| \to +\infty$ is presented by $MG_\alpha$.
This idea is developed from pointwise-estimates for Navier-Stokes flow via Miyakawa \cite{Mykw00}, and Miyakawa and Schonbek \cite{Mykw-Schnbk}, and firstly applied by Brandolese \cite{Brndls}, and Brandolese and Vigneron \cite{Brndls-Vgnrn}.
A main goal of this paper is to show the uniform estimate of the spatial decay of the solution.
Specifically, we provide the similar estimate as \eqref{L2} in $L^\infty (\mathbb{R}^2)$.
For solutions to the fractional diffusion equation $\partial_t u + (-\Delta)^{\alpha/2} u = \nabla \cdot f(u)$ with $1 < \alpha < 2$ and some suitable $f$, general theory of spatial decay is given by Brandolese and Karch \cite{Brndls-Krch}.
This theory is based on the $L^p$-$L^q$ estimates for $e^{-t(-\Delta)^{\alpha/2}}$ and available for \eqref{qg} in the subcritical case $1 < \alpha < 2$ since \eqref{qg} is parabolic in this case.
However, since the nonlinearity balances to the dissipation in $\alpha = 1$, the general theory does not work in the case that $0 < \alpha \le 1$.
More precisely, it is diffucult to estimate the nonlinear term in the integral equation $\int_0^t \nabla e^{-(t-s)(-\Delta)^{\alpha/2}} \cdot f(u) ds$ in usual way, since $\nabla e^{-t(-\Delta)^{\alpha/2}}$ is not integrable near $t = 0$, which requires to estimate $f(u)$ in the weighted Sobolev spaces with some positive differential order.
In particular, we estimate $\| |x|^2 \nabla \theta (t) \|_{L^p (\mathbb{R}^2)}$ by using the energy method (cf.\cite{Y-S-NA}).
Furthermore we prepare some uniform decay properties for $G_\alpha$ (see Lemmas \ref{prop-wt} and \ref{lem1}).
We note that the Moser-Nash iteration method is not used, which is usually employed in order to obtain $L^\infty$ estimate from $L^p$ (e.g.\cite{E-Z,K-K,Og-Y}).
In \cite{Y-S18aX}, the estimate \eqref{L2} is derived by the energy method.
The proof of the uniform estimate \eqref{astn} in the main theorem is based on the $L^p$-$L^q$ argument.
We can express the spatial decay of the solution both in the critical and the supercritical cases.

\begin{theorem}\label{thm}
Let $0 < \alpha \le 1,~ p > 2/\alpha,~ (1+|x|) \theta_0 \in L^1 (\mathbb{R}^2),~ |x|^{3+\alpha} \theta_0 \in L^\infty (\mathbb{R}^2),~ \theta_0 \in H^3 (\mathbb{R}^2)$ and $|x|^2 \nabla \theta_0 \in L^p (\mathbb{R}^2)$.
Assume that $\| \theta_0 \|_{H^3 (\mathbb{R}^2)}$ is sufficiently small, and the solution $\theta$ fulfills \eqref{exist} and \eqref{mass}.
Then
\begin{equation}\label{astn}
	\bigl\| |x|^{3+\alpha} \bigl( \theta (t) - MG_\alpha (t) \bigr) \bigr\|_{L^\infty (\mathbb{R}^2)}
	\le
	C (1+t),
\end{equation}
where $M = \int_{\mathbb{R}^2} \theta_0 (x) dx$.
\end{theorem}
We emphasize that $\| |x|^{3+\alpha} G_\alpha (t) \|_{L^\infty (\mathbb{R}^2)} = + \infty$.
Thus Theorem \ref{thm} states that $\theta$ and $MG_\alpha$ are canceled uniformly in far field.
We remark that the assertions in this theorem is sharp in time.
Indeed, \eqref{astn} for $G_\alpha (t) * \theta_0$ instead of $\theta$ is fulfilled (see Lemma \ref{lem1} in Section 2), and the scaling property of $G_\alpha$ guarantees the sharpness of those estimates.
Namely, from the mean value theorem, we expect that the decay-rate of the top term of $G_\alpha (t) * \theta_0 - M G_\alpha (t)$ is given by one of $\nabla G_\alpha (t)$, and $\| |x|^{3+\alpha} \nabla G_\alpha (t) \|_{L^\infty (\mathbb{R}^2)} = t \| |x|^{3+\alpha} \nabla G_\alpha (1) \|_{L^\infty (\mathbb{R}^2)}$ and $\| |x|^{3+\alpha} \nabla G_\alpha (1) \|_{L^\infty (\mathbb{R}^2)} < +\infty$.
The details are in the proof of Lemma \ref{lem1}.
Furthermore \eqref{astn} is optimal also in $x$ since $\| |x|^{3+\alpha+\varepsilon} \nabla G_\alpha (t) \|_{L^\infty (\mathbb{R}^2)} = + \infty$ for any $\varepsilon > 0$.
A coupling of this theorem and the property of $G_\alpha$ provides the obvious decay as follows.
\begin{cor}\label{cor}
Let $\varrho = \varrho(x) > 0$ be radially symmetric and monotone increasing in $|x|$, and satisfy that $\varrho (x) \to + \infty$ as $|x| \to +\infty$.
Let $0 < \alpha \le 1,~ p > 2/\alpha,~ \theta_0 \in L^1 (\mathbb{R}^2) \cap H^3 (\mathbb{R}^2),~ |x|^{2+\alpha} \varrho\, \theta_0 \in L^\infty (\mathbb{R}^2)$ and $|x|^2 \nabla \theta_0 \in L^p (\mathbb{R}^2)$.
Assume that the solution $\theta$ fulfills \eqref{exist}, \eqref{mass}, and \eqref{drbdd} for $\sigma = 3$.
Then, for any fixed $t>0$,
\[
	|x|^{2+\alpha} \theta (t,x)
	\to C_\alpha M t
\]
as $|x| \to +\infty$, where $M = \int_{\mathbb{R}^2} \theta_0 (x) dx$ and $C_\alpha = \alpha 2^{\alpha- 1} \pi^{-2} \sin \tfrac{\alpha \pi}2 \Gamma (1+\tfrac{\alpha}2) \Gamma (\tfrac\alpha{2})$.
\end{cor}
We should remark that $\theta$ on Theorem \ref{thm} and Corollary \ref{cor} fulfill \eqref{wttheta} for any $q > 2/\alpha$.
Indeed, if $|x|^2 \theta_0 \in L^q (\mathbb{R}^2)$, then Proposition \ref{prop-wt} in Section 2 yields \eqref{wttheta}, and we confirm that
\[
	\bigl\| |x|^2 \theta_0 \bigr\|_{L^q (\mathbb{R}^2)}
	\le
	\bigl\| (1+ |x|)^{-\alpha} \bigr\|_{L^q (\mathbb{R}^2)} \bigl\| |x|^2 (1+|x|)^{\alpha} \theta_0 \bigr\|_{L^\infty (\mathbb{R}^2)}
	< + \infty.
\]

\noindent
{\bf Notation.}
The Fourier transform and its inverse are defined by $\mathcal{F} [\varphi] (\xi) = (2\pi)^{-1} \int_{\mathbb{R}^2} e^{-ix\cdot\xi} \varphi (x) dx$ and $\mathcal{F}^{-1} [\varphi] (x) = (2\pi)^{-1} \int_{\mathbb{R}^2} e^{ix\cdot\xi} \varphi (\xi) d\xi$, where $i = \sqrt{-1}$.
The derivations are abbreviated by $\partial_t = \partial / \partial t,~ \partial_j = \partial / \partial x_j$ for $j = 1,2,~ \nabla = (\partial_1,\partial_2),~ \nabla^\bot = (-\partial_2,\partial_1)$ and $\Delta = \partial_1^2 + \partial_2^2$.
The fractional Laplacian and its inverse, and the Riesz transform are defined by $(-\Delta)^{\alpha/2} \varphi = \mathcal{F}^{-1} [|\xi|^\alpha \mathcal{F} [\varphi]],~ (-\Delta)^{-\sigma/2} \varphi = \mathcal{F}^{-1} [|\xi|^{-\sigma} \mathcal{F} [\varphi]]$ for $0 < \sigma < 2$, and $R_j \varphi = \partial_j (-\Delta)^{-1/2} \varphi = \mathcal{F}^{-1} [i\xi_j |\xi|^{-1} \mathcal{F} [\varphi]]$ for $j = 1,2$, respectively.
The H\"older conjugate of $1 \le p \le \infty$ is denoted by $p'$, i.e., $\frac1p + \frac1{p'} = 1$.
For $\beta = (\beta_1,\beta_2) \in \mathbb{Z}_+^2 = (\mathbb{N} \cup \{ 0 \})^2,~ |\beta| = \beta_1 + \beta_2$.
For some operators $A$ and $B,~ [A,B] = AB - BA$.
Various positive constants and suitable fuctions are denoted by $C$ and $\varphi$, respectively.

\section{Preliminaries}
The Duhamel principle yields that
\begin{equation}\label{MS}
	\theta (t)
	=
	G_\alpha (t) * \theta_0
	-
	\int_0^t
		\nabla G_\alpha (t-s) * (\theta \nabla^\bot \psi) (s)
	ds,
\end{equation}
where $G_\alpha (t) = (2\pi)^{-1} \mathcal{F}^{-1} [e^{-t|\xi|^\alpha}]$.
It is well-known that
\begin{equation}\label{scG}
	G_\alpha (t,x) = t^{-2/\alpha} G_\alpha (1,t^{-1/\alpha} x)
\end{equation}
for $(t,x) \in (0,+\infty) \times \mathbb{R}^2$, and
\begin{equation}\label{decayG}
	\bigl| \nabla^\beta G_\alpha (1,x) \bigr|
	\le
	C_\beta (1+|x|^2)^{-1-\frac\alpha{2}-\frac{|\beta|}2}
\end{equation}
for $\beta \in \mathbb{Z}_+^2$ and $x \in \mathbb{R}^2$ (see\cite{Klkltsv,Y-S18aX}).
The spatial decay of $G_\alpha$ is published as the following (cf.\cite{Blmntr-Gtr}):
\begin{equation}\label{B-G}
	|x|^{2+\alpha} G_\alpha (t,x) \to C_\alpha t
\end{equation}
as $|x| \to +\infty$, where $C_\alpha$ is introduced in Corollary \ref{cor}.
The following lemma plays a crucial role in the energy estimates.
\begin{lemma}[Stroock-Varopoulos inequality \cite{Crdb-Crdb,J05,L-S}]\label{lemC-C}
	Let $0 \le \alpha \le 2,~ q \ge 2$ and $f \in W^{\alpha,q} (\mathbb{R}^2)$.
	Then
	\[
		\int_{\mathbb{R}^2}
			|f|^{q-2} f (-\Delta)^{\alpha/2} f
		dx
		\ge
		\frac2q \int_{\mathbb{R}^2}
			\left| (-\Delta)^{\alpha/4} (|f|^{q/2}) \right|^2
		dx
	\]
	holds.
\end{lemma}
We need the following inequalities of Sobolev type.
\begin{lemma}[Hardy-Littlewood-Sobolev's inequality \cite{S,Z}]\label{HLS}~
	Let $0 < \sigma < 2,~ 1 < p < \frac{2}{\sigma}$ and $\frac1{p_*} = \frac1p - \frac{\sigma}2$.
	Then there exists a positive constant $C$ such that
	\[
		\bigl\| (-\Delta)^{-\sigma/2} \varphi \bigr\|_{L^{p_*} (\mathbb{R}^2)}
		\le
		C \bigl\| \varphi \bigr\|_{L^p (\mathbb{R}^2)}
	\]
	for any $\varphi \in L^p (\mathbb{R}^2)$.
\end{lemma}
\begin{lemma}[Gagliardo-Nirenberg inequality \cite{H-Y-Z,K-PP,M-N-S-S}]\label{GN}
	Let $0 < \sigma < s < 2,~ 1 < p_1, p_2 < \infty$ and $\frac1p = (1-\frac\sigma{s}) \frac1{p_1} + \frac\sigma{s} \frac1{p_2}$.
	Then
	\[
		\bigl\| (-\Delta)^{\sigma/2} \varphi \bigr\|_{L^p (\mathbb{R}^2)}
		\le
		C \bigl\| \varphi \bigr\|_{L^{p_1} (\mathbb{R}^2)}^{1-\frac\sigma{s}}
		\bigl\| (-\Delta)^{s/2} \varphi \bigr\|_{L^{p_2} (\mathbb{R}^2)}^{\frac\sigma{s}}
	\]
	holds.
\end{lemma}
To care the Riesz transforms, we call the following H\"ormander-Mikhlin type estimate.
\begin{lemma}[H\"ormander-Mikhlin inequality\,\cite{Hrmndr,Mkhln,S-S}]\label{lem-HM}
Let $N \in \mathbb{Z}_+,~ 0 < \mu \le 1$ and $\lambda = N + \mu - 2$.
Assume that $\varphi \in C^\infty (\mathbb{R}^2\backslash \{ 0 \})$ satisfies the following conditions:
\begin{itemize}
\item
	$\nabla^\gamma \varphi \in L^1 (\mathbb{R}^2)$ for any $\gamma \in \mathbb{Z}_+^2$ with $|\gamma| \le N$;
\item
	$|\nabla^\gamma \varphi (\xi)| \le C_\gamma |\xi|^{\lambda - |\gamma|}$ for $\xi \neq 0$ and $\gamma \in \mathbb{Z}_+^2$ with $|\gamma| \le N+2$.
\end{itemize}
Then
\[
	\sup_{x \neq 0} \bigl( |x|^{2+\lambda} \bigl| \mathcal{F}^{-1} [\varphi] (x) \bigr| \bigr)
	< + \infty
\]
holds.
\end{lemma}
For the details of this Lemma, see \cite{S-S}.
The following proposition is confirmed in \cite{Crdb-Crdb,J04,Y-S18aX}.
\begin{proposition}\label{tr}
Let $\sigma > 2,~ \theta_0 \in H^\sigma (\mathbb{R}^2)$ and $\| \theta_0 \|_{H^\sigma (\mathbb{R}^2)}$ be small.
Assume that the solution $\theta$ satisfies \eqref{exist} and \eqref{mass}.
Then \eqref{drbdd} holds.
\end{proposition}
The authors proved the following proposition in \cite{Y-S18aX}.
\begin{proposition}\label{prop-wt}
Let $q > 2/\alpha$ and $|x|^2 \theta_0 \in L^q (\mathbb{R}^2)$.
Assume that the solution $\theta$ of \eqref{qg} satiefies \eqref{exist} and \eqref{mass}.
Then \eqref{wttheta} holds.
\end{proposition}
The term of initial-data on \eqref{MS} satisfies the following lemma.
\begin{lemma}\label{lem1}
Let $(1+|x|) \theta_0 \in L^1 (\mathbb{R}^2)$ and $|x|^{3+\alpha} \theta_0 \in L^\infty (\mathbb{R}^2)$.
Then
\[
	\bigl\| |x|^{3+\alpha} \bigl( G_\alpha (t) * \theta_0 - M G_\alpha (t) \bigr) \bigr\|_{L^\infty (\mathbb{R}^2)}
	\le C \bigl( t \bigl\| |x| \theta_0 \bigr\|_{L^1 (\mathbb{R}^2)} + \bigl\| |x|^{3+\alpha} \theta_0 \bigr\|_{L^\infty (\mathbb{R}^2)} \bigr)
\]
for $t > 0$, where $M = \int_{\mathbb{R}^2} \theta_0 (x) dx$.
\end{lemma}
\begin{proof}
The mean value theorem gives that
\begin{equation}\label{meanG}
\begin{split}
	G_\alpha (t) * \theta_0 - M G_\alpha (t)
	&=
	\int_{|y| > |x|/2} \bigl( G_\alpha (t,x-y) - G_\alpha (t,x) \bigr) \theta_0 (y) dy\\
	&+
	\int_{|y| \le |x|/2} \int_0^1 (-y\cdot\nabla) G_\alpha (t,x-\lambda y) \theta_0 (y) d\lambda dy.
\end{split}
\end{equation}
The first term fulfills that
\[
\begin{split}
	&\biggl| |x|^{3+\alpha} \int_{|y| > |x|/2} \bigl( G_\alpha (t,x-y) - G_\alpha (t,x) \bigr) \theta_0 (y) dy \biggr|\\
	&\le
	C \bigl\| G_\alpha (t) \bigr\|_{L^1 (\mathbb{R}^2)} \bigl\| |x|^{3+\alpha} \theta_0 \bigr\|_{L^\infty (\mathbb{R}^2)}
	+
	C \bigl\| |x|^{2+\alpha} G_\alpha (t) \bigr\|_{L^\infty (\mathbb{R}^2)} \bigl\| |x| \theta_0 \bigr\|_{L^1 (\mathbb{R}^2)}\\
	&\le
	 C \bigl( t \bigl\| |x| \theta_0 \bigr\|_{L^1 (\mathbb{R}^2)} + \bigl\| |x|^{3+\alpha} \theta_0 \bigr\|_{L^\infty (\mathbb{R}^2)} \bigr).
\end{split}
\]
From \eqref{scG} and \eqref{decayG}, for $|y| \le |x|/2$ and $0 < \lambda < 1$, we see that $|x| \le 2|x-\lambda y|$ and then
\[
\begin{split}
	|x|^{3+\alpha} \bigl| \nabla G_\alpha (t,x-\lambda y) \bigr|
	&=
	t^{-3/\alpha} |x|^{3+\alpha} \bigl| \nabla G_\alpha (1,t^{-1/\alpha}(x-\lambda y)) \bigr|
	\le
	C t.
\end{split}
\]
Thus, we have that
\[
	\biggl| |x|^{3+\alpha} \int_{|y| \le |x|/2} \int_0^1 (-y\cdot\nabla) G_\alpha (t,x-\lambda y) \theta_0 (y) d\lambda dy \biggr|
	\le
	C t \bigl\| |x| \theta_0 \bigr\|_{L^1 (\mathbb{R}^2)}
\]
and conclude the proof.
\end{proof}
\begin{lemma}\label{lem2}
Let $\theta_0 \in L^1 (\mathbb{R}^2)$ and $|x|^{2+\alpha} \varrho\, \theta_0 \in L^\infty (\mathbb{R}^2)$, where $\varrho = \varrho (x)$ is defined in Corollary \ref{cor}.
Then, for any $t > 0$,
\[
	|x|^{2+\alpha} \bigl( G_\alpha (t) * \theta_0 - M G_\alpha (t) \bigr)
	\to 0
\]
as $|x| \to +\infty$, where $M = \int_{\mathbb{R}^2} \theta_0 (x) dx$.
\end{lemma}
\begin{proof}
We use \eqref{meanG}.
For the first term, we see that
\[
\begin{split}
	&|x|^{2+\alpha} \biggl| \int_{|y| > |x|/2} \bigl( G_\alpha (t,x-y) - G_\alpha (t,x) \bigr) \theta_0 (y) dy \biggr|\\
	&\le
	C \varrho (\tfrac{x}2)^{-1} \int_{|y| > |x|/2} \bigl| G_\alpha (t,x-y) \bigr| \bigl| |y|^{2+\alpha} \varrho(y) \theta_0 (y) \bigr| dy
	+ C \int_{|y| > |x|/2} \bigl| |x|^{2+\alpha} G_\alpha (t,x) \bigr| \bigl| \theta_0 (y) \bigr| dy\\
	&\le
	C \varrho (\tfrac{x}2)^{-1} \bigl\| |x|^{2+\alpha} \varrho \theta_0 \bigr\|_{L^\infty (\mathbb{R}^2)}
	+
	C \int_{|y| > |x|/2} \bigl| \theta_0 (y) \bigr| dy
	\to 0
\end{split}
\]
as $|x| \to +\infty$.
For the second term on \eqref{meanG}, we choose sufficiently small $\varepsilon > 0$ and $\frac2{1+\alpha} < r < \frac2{1+\varepsilon}$, then we have by the similar argument as in the proof of Lemma \ref{lem1} that
\[
\begin{split}
	&|x|^{2+\alpha} \biggl| \int_{|y| \le |x|/2} \int_0^1 (-y\cdot\nabla) G_\alpha (t,x-\lambda y) \theta_0 (y) d\lambda dy \biggr|\\
	&\le
	C |x|^{-\varepsilon} \int_{|y| \le |x|/2} \int_0^1 |x-\lambda y|^{2+\alpha+\varepsilon} \bigl| \nabla G_\alpha (t,x-\lambda y) \bigr|
	\bigl| y \theta_0 (y) \bigr| d\lambda dy\\
	&\le
	C |x|^{-\varepsilon} t^{-\frac2{\alpha r} + \frac1\alpha + 1 + \frac\varepsilon\alpha} \bigl\| x \theta_0 \bigr\|_{L^r (\mathbb{R}^2)}
	\to 0
\end{split}
\]
as $|x| \to +\infty$.
Here we remark that $-\frac2{\alpha r} + \frac1\alpha + 1 + \frac\varepsilon\alpha > 0$ and
\[
	\bigl\| x \theta_0 \bigr\|_{L^r (\mathbb{R}^2)}
	\le
	 \bigl\| (1+|x|)^{-1-\alpha} \bigr\|_{L^r (\mathbb{R}^2)} \bigl\| |x| (1+|x|)^{1+\alpha} \theta_0 \bigr\|_{L^\infty (\mathbb{R}^2)}
	< + \infty.
\]
Hence we complete the proof.
\end{proof}
\section{Proof of main theorems}
To show our main assertions, we prepare the estimate for the nonlinear effect.
We denote the nonlinear term on \eqref{MS} by $v$, i.e.,
\begin{equation}\label{def-v}
	v (t) = 
	-
	\int_0^t
		\nabla G_\alpha (t-s) * (\theta \nabla^\bot \psi) (s)
	ds.
\end{equation}
Then decay-rate of $v$ as $|x| \to +\infty$ is published as follows.
\begin{proposition}\label{prop-v2}
Let $0 < \alpha \le 1,~ p > 2/\alpha,~ \theta_0 \in L^1 (\mathbb{R}^2) \cap H^3 (\mathbb{R}^2),~ |x|^{2+\alpha} \theta_0 \in L^\infty (\mathbb{R}^2)$ and $|x|^2 \nabla \theta_0 \in L^p (\mathbb{R}^2)$.
Assume that the solution $\theta$ fulfills \eqref{mass} and \eqref{drbdd} for $\sigma = 3$.
Then $v$ defined by \eqref{def-v} satisfies that
$\| |x|^{3+\alpha} v(t) \|_{L^\infty (\mathbb{R}^2)} \le C (1+t)$.
\end{proposition}
\begin{proof}
From the definition, we see for $j = 1,2$ that
\[
\begin{split}
	x_j v &= - \int_0^t (x_j \nabla G_\alpha) (t-s) * (\theta \nabla^\bot\psi) (s) ds\\
	&- \int_0^t G_\alpha (t-s) * (\theta(\nabla^\bot\psi)_j) (s) ds
	- \int_0^t G_\alpha (t-s) * (x_j \nabla\theta\cdot\nabla^\bot\psi) (s) ds,
\end{split}
\]
where $(\nabla^\bot\psi)_j$ is the $j$-th component of $\nabla^\bot\psi$.
Hence
\begin{equation}\label{bss}
\begin{split}
	&\bigl\| |x|^{2+\alpha} x_j v \bigr\|_{L^\infty (\mathbb{R}^2)}
	\le
	C \int_0^t
		\bigl\| |x|^{2+\alpha} x_j \nabla G_\alpha (t-s) \bigr\|_{L^\infty (\mathbb{R}^2)}
		\bigl\| \theta \nabla^\bot \psi \bigr\|_{L^1 (\mathbb{R}^2)}
	ds\\
	&+
	C \int_0^t \int_{\mathbb{R}^2}
		\bigl|  (x_j-y_j) \nabla G_\alpha (t-s,x-y) \bigr|
		\bigl| |y|^{2+\alpha} (\theta \nabla^\bot \psi)(s,y) \bigr|
	dyds\\
	&+
	\int_0^t
		\bigl\| |x|^{2+\alpha} G_\alpha (t-s) \bigr\|_{L^\infty (\mathbb{R}^2)}
		\bigl\| \theta (\nabla^\bot\psi)_j (s) \bigr\|_{L^1 (\mathbb{R}^2)}
	ds\\
	&+
	\int_0^t \int_{\mathbb{R}^2}
		\bigl| G_\alpha (t-s,x-y) \bigr|
		\bigl| |y|^{2+\alpha} (\theta (\nabla^\bot\psi)_j) (s,y) \bigr|
	dyds\\
	&+
	\int_0^t
		\bigl\| |x|^{2+\alpha} G_\alpha (t-s) \bigr\|_{L^\infty (\mathbb{R}^2)}
		\bigl\| x_j (\nabla \theta \cdot \nabla^\bot\psi) (s) \bigr\|_{L^1 (\mathbb{R}^2)}
	ds\\
	&+
	\int_0^t \int_{\mathbb{R}^2}
		\bigl| G_\alpha (t-s,x-y) \bigr|
		\bigl| |y|^{2+\alpha} y_j (\nabla \theta \cdot \nabla^\bot\psi) (s,y) \bigr|
	dyds.
\end{split}
\end{equation}
From \eqref{mass}-\eqref{wttheta}, \eqref{scG} and \eqref{decayG}, we see that the first and the third terms are bounded by $C (1+t)$.
We estimate the second, the fourth and the fifth terms later.
To care the last term, we prepare the estimate for $\| |x|^2 \nabla \theta \|_{L^p (\mathbb{R}^2)}$.
We derivate the first equality on \eqref{qg} in $x_k$ and multiply $|x|^{2p} |\partial_k \theta|^{p-2} \partial_k \theta$.
Then, by integrating it in $(0,t) \times \mathbb{R}^2$ and employing Lemma \ref{lemC-C}, we see for the second and the last terms that
\[
\begin{split}
	&\int_{\mathbb{R}^2}
		|x|^{2p} |\partial_k \theta|^{p-2} \partial_k \theta (-\Delta)^{\alpha/2} \partial_k \theta
	dx\\
	&=
	\int_{\mathbb{R}^2}
		\bigl| |x|^2 \partial_k \theta \bigr|^{p-2} |x|^2 \partial_k \theta (-\Delta)^{\alpha/2} (|x|^2 \partial_k \theta)
	dx
	-
	\int_{\mathbb{R}^2}
		\bigl| |x|^2 \partial_k \theta \bigr|^{p-2} |x|^2 \partial_k \theta \bigl[ (-\Delta)^{\alpha/2}, |x|^2 \bigr] \partial_k \theta
	dx\\
	&\ge
	\frac2p \bigl\| (-\Delta)^{\alpha/4} \bigl( \bigl| |x|^2 \partial_k \theta \bigr|^{p/2} \bigr) \bigr\|_{L^2 (\mathbb{R}^2)}^2
	-
	\int_{\mathbb{R}^2}
		\bigl| |x|^2 \partial_k \theta \bigr|^{p-2} |x|^2 \partial_k \theta \bigl[ (-\Delta)^{\alpha/2}, |x|^2 \bigr] \partial_k \theta
	dx
\end{split}
\]
and
\[
\begin{split}
	&\int_{\mathbb{R}^2}
		|x|^{2p} |\partial_k \theta|^{p-2} \partial_k \theta \partial_k \nabla \cdot (\theta \nabla^\bot\psi)
	dx\\
	&=
	-2p \int_{\mathbb{R}^2}
		\bigl| |x|^2 \partial_k \theta \bigr|^{p-2} |x|^2 \partial_k \theta x_k \nabla\theta \cdot \nabla^\bot \psi
	dx
	-(p-1) \int_{\mathbb{R}^2}
		\bigl| |x|^2 \partial_k \theta \bigr|^{p-2} \partial_k^2 \theta |x|^4 \nabla\theta\cdot\nabla^\bot\psi
	dx.
\end{split}
\]
Here we used the relation $\nabla \cdot (\theta \nabla^\bot\psi) = \nabla \theta \cdot \nabla^\bot \psi$.
Therefore we have that
\begin{equation}\label{bsg}
\begin{split}
	&\frac1p \sum_{k=1}^2 \bigl\| |x|^2 \partial_k \theta \bigr\|_{L^p (\mathbb{R}^2)}^p
	+
	\frac2p \sum_{k=1}^2 \int_0^t \bigl\| (-\Delta)^{\alpha/4} \bigl( \bigl| |x|^2 \partial_k \theta \bigr|^{p/2} \bigr) \bigr\|_{L^2 (\mathbb{R}^2)}^2 ds\\
	&\le
	\frac1p \sum_{k=1}^2 \bigl\| |x|^2 \partial_k \theta_0 \bigr\|_{L^p (\mathbb{R}^2)}^p
	+
	2p \sum_{k=1}^2 \int_0^t \int_{\mathbb{R}^2}
		\bigl| |x|^2 \partial_k \theta \bigr|^{p-2} |x|^2 \partial_k \theta x_k \nabla\theta \cdot \nabla^\bot \psi
	dxds\\
	&+
	(p-1) \sum_{k=1}^2 \int_0^t \int_{\mathbb{R}^2}
		\bigl| |x|^2 \partial_k \theta \bigr|^{p-2} \partial_k^2 \theta |x|^4 \nabla\theta\cdot\nabla^\bot\psi
	dxds\\
	&+
	\sum_{k=1}^2 \int_0^t \int_{\mathbb{R}^2}
		\bigl| |x|^2 \partial_k \theta \bigr|^{p-2} |x|^2 \partial_k \theta \bigl[ (-\Delta)^{\alpha/2}, |x|^2 \bigr] \partial_k \theta
	dxds.
\end{split}
\end{equation}
From \eqref{MS} and Hausdorf-Young's inequality, we see that
\[
\begin{split}
	\bigl\| |x|^2 R_l \theta \bigr\|_{L^\infty (\mathbb{R}^2)}
	&\le
	\bigl\| |x|^2 (R_l G_\alpha (t) * \theta_0) \bigr\|_{L^\infty (\mathbb{R}^2)}
	+
	\biggl\| |x|^2 \int_0^t R_l G_\alpha (t-s) * (\nabla\theta \cdot \nabla^\bot \psi) (s) ds \biggr\|_{L^\infty (\mathbb{R}^2)}\\
	&\le
	C \Bigl( 1 + \bigl\| \theta_0 \bigr\|_{H^2 (\mathbb{R}^2)} \bigl\| |x|^2 \nabla^\bot \psi \bigr\|_{L^\infty ((0,\infty)\times\mathbb{R}^2)} \Bigr).
\end{split}
\]
Here we used the scaling property $\| |x|^2 R_l G_\alpha (t) \|_{L^\infty (\mathbb{R}^2)} = \| |x|^2 R_l G_\alpha (1) \|_{L^\infty (\mathbb{R}^2)}$.
We remark that Lemma \ref{lem-HM} guarantees that $|x|^2 R_l G_\alpha (1) \in L^\infty (\mathbb{R}^2)$.
The relation $\nabla^\bot \psi = (-R_2 \theta, R_1 \theta)$ and the smallness of $\theta_0$ conclude that $\| |x|^2 \nabla^\bot \psi \|_{L^\infty (\mathbb{R}^2)} \le C$.
Hence
for the second term on \eqref{bsg} and some small $\delta > 0$, we have that
\begin{equation}\label{bsg2}
\begin{split}
	&\sum_{k=1}^2 \biggl| \int_0^t \int_{\mathbb{R}^2}
		\bigl| |x|^2 \partial_k \theta \bigr|^{p-2} |x|^2 \partial_k \theta x_k \nabla\theta \cdot \nabla^\bot \psi
	dxds \biggr|\\
	&\le
	C \sum_{k=1}^2 \int_0^t
		\bigl\| |x|^2 \partial_k \theta \bigr\|_{L^p (\mathbb{R}^2)}^{p-1}
		\bigl\| |x|^2  \nabla^\bot \psi \bigr\|_{L^\infty (\mathbb{R}^2)}^{1/2}
		\bigl\| |\nabla^\bot\psi|^{1/2} \nabla\theta \bigr\|_{L^p (\mathbb{R}^2)}
	ds\\
	&\le
	C \sum_{k=1}^2 \bigl\| |x|^2 \partial_k \theta \bigr\|_{L^\infty (0,t: L^p (\mathbb{R}^2))}^{p-1}
	 \int_0^t
		(1+s)^{-\frac2\alpha (1-\frac1p)-\frac2\alpha}
	ds
	\le
	C_\delta + \delta \sum_{k=1}^2 \bigl\| |x|^2 \partial_k \theta \bigr\|_{L^\infty (0,t: L^p (\mathbb{R}^2))}^p.
\end{split}
\end{equation}
For the third term on \eqref{bsg}, we see for $\frac1r = \frac1p + \frac\alpha{2} - \frac\alpha{2p}$ that
\[
\begin{split}
	&\sum_{k=1}^2 \biggl| \int_0^t \int_{\mathbb{R}^2}
		\bigl| |x|^2 \partial_k \theta \bigr|^{p-2} \partial_k^2 \theta |x|^4 \nabla\theta\cdot\nabla^\bot\psi
	dxds \biggr|\\
	&\le
	C \sum_{k,l=1}^2\int_0^t
		\bigl\| (-\Delta)^{\alpha/4} \bigl( \bigl| |x|^2 \partial_k \theta \bigr|^{p/2} \bigr) \bigr\|_{L^2 (\mathbb{R}^2)}^{2/p'}
		\bigl\| \partial_l^2 \theta \bigr\|_{L^r (\mathbb{R}^2)}
	ds.
\end{split}
\]
Here we used the Sobolev inequality $\| \varphi \|_{L^{2_*} (\mathbb{R}^2)} \le C \| (-\Delta)^{\alpha/4} \varphi \|_{L^2 (\mathbb{R}^2)}$ for $\frac1{2_*} = \frac12 - \frac\alpha{4}$.
Lemma \ref{GN} with Proposition \ref{tr} guarantees that $\partial_l^2 \theta \in L^r (\mathbb{R}^2)$ and $\| \partial_l^2 \theta \|_{L^r (\mathbb{R}^2)} \le C (1+t)^{-\frac2\alpha (1-\frac1r) - \frac2\alpha}$.
Hence
\begin{equation}\label{bsg3}
\begin{split}
	&\sum_{k=1}^2 \biggl| \int_0^t \int_{\mathbb{R}^2}
		\bigl| |x|^2 \partial_k \theta \bigr|^{p-2} \partial_k^2 \theta |x|^4 \nabla\theta\cdot\nabla^\bot\psi
	dxds \biggr|\\
	&\le
	C_\delta \int_0^t (1+s)^{-\frac2\alpha (p-1) - \frac{2p}\alpha -1 + p} ds
	+
	\delta \sum_{k=1}^2\int_0^t
		\bigl\| (-\Delta)^{\alpha/4} \bigl( \bigl| |x|^2 \partial_k \theta \bigr|^{p/2} \bigr) \bigr\|_{L^2 (\mathbb{R}^2)}^2
	ds\\
	&\le
	C_\delta
	+
	\delta \sum_{k=1}^2\int_0^t
		\bigl\| (-\Delta)^{\alpha/4} \bigl( \bigl| |x|^2 \partial_k \theta \bigr|^{p/2} \bigr) \bigr\|_{L^2 (\mathbb{R}^2)}^2
	ds.
\end{split}
\end{equation}
Calculus on the Fourier symbol yield that $[ (-\Delta)^{\alpha/2}, |x|^2 ] \partial_k \theta = \alpha^2 (-\Delta)^{\frac{\alpha-2}2} \partial_k \theta - 2\alpha \nabla (-\Delta)^{\frac{\alpha-2}2} \cdot (x \partial_k \theta)$.
Thus, for $\frac1r = \frac1p + \frac{1-\alpha}2$,
\[
\begin{split}
	&\sum_{k=1}^2 \biggl|
		\int_0^t \int_{\mathbb{R}^2}
		\bigl| |x|^2 \partial_k \theta \bigr|^{p-2} |x|^2 \partial_k \theta \bigl[ (-\Delta)^{\alpha/2}, |x|^2 \bigr] \partial_k \theta
	dxds \biggr|\\
	&\le
	C \sum_{k=1}^2 \bigl\| |x|^2 \partial_k \theta \bigr\|_{L^\infty (0,t: L^p (\mathbb{R}^2))}^{p-1}
	\int_0^t
		\Bigl( \bigl\| (-\Delta)^{\frac{\alpha-1}2} (x \partial_k \theta) \bigr\|_{L^p (\mathbb{R}^2)}
		+ \bigl\| (-\Delta)^{\frac{\alpha-2}2} \partial_k \theta \bigr\|_{L^p (\mathbb{R}^2)} \Bigr)
	ds\\
	&\le
	C \sum_{k=1}^2 \bigl\| |x|^2 \partial_k \theta \bigr\|_{L^\infty (0,t: L^p (\mathbb{R}^2))}^{p-1}
	\int_0^t
		\Bigl( \bigl\| x \partial_k \theta \bigr\|_{L^r (\mathbb{R}^2)} + \bigl\| \theta \bigr\|_{L^r (\mathbb{R}^2)} \Bigr)
	ds.
\end{split}
\]
Since $\frac{pr}{2p-r} > 1$, we have from the Sobolev inequality with \eqref{wttheta} that
\[
	\bigl\| x \partial_k \theta \bigr\|_{L^r (\mathbb{R}^2)}
	\le
	C \bigl\| \partial_k \theta \bigr\|_{L^{\frac{pr}{2p-r}} (\mathbb{R}^2)}^{1/2} \bigl\| |x|^2 \partial_k \theta \bigr\|_{L^p (\mathbb{R})}^{1/2}
	\le
	C (1+t)^{-\frac1{2\alpha}-1+\frac1{\alpha p}} \bigl\| |x|^2 \partial_k \theta \bigr\|_{L^p (\mathbb{R})}^{1/2}.
\]
Therefore
\begin{equation}\label{bsg4}
\begin{split}
	&\sum_{k=1}^2 \biggl|
		\int_0^t \int_{\mathbb{R}^2}
		\bigl| |x|^2 \partial_k \theta \bigr|^{p-2} |x|^2 \partial_k \theta \bigl[ (-\Delta)^{\alpha/2}, |x|^2 \bigr] \partial_k \theta
	dxds \biggr|\\
	&\le
	C \sum_{k=1}^2 \bigl\| |x|^2 \partial_k \theta \bigr\|_{L^\infty (0,t: L^p (\mathbb{R}^2))}^{p-\frac12}
	\int_0^t (1+s)^{-\frac1{2\alpha}-1+\frac1{\alpha p}} ds\\
	&+
	C \sum_{k=1}^2 \bigl\| |x|^2 \partial_k \theta \bigr\|_{L^\infty (0,t: L^p (\mathbb{R}^2))}^{p-1}
	\int_0^t (1+s)^{-\frac2\alpha (1-\frac1p)-1+\frac1\alpha} ds
	\le
	C_\delta + \delta  \sum_{k=1}^2 \bigl\| |x|^2 \partial_k \theta \bigr\|_{L^\infty (0,t: L^p (\mathbb{R}^2))}^p.
\end{split}
\end{equation}
Consequently, applying \eqref{bsg2}-\eqref{bsg4} to \eqref{bsg},
\begin{equation}\label{tokyu}
	 \bigl\| |x|^2 \nabla \theta \bigr\|_{L^\infty (0,\infty: L^p (\mathbb{R}^2))}^p
	 +
	 \int_0^\infty \bigl\| (-\Delta)^{\alpha/4} \bigl( \bigl| |x|^2 \nabla \theta \bigr|^{p/2} \bigr) \bigr\|_{L^2 (\mathbb{R}^2)}^2 ds
	 < + \infty.
\end{equation}
We back to the estimate for the last term on \eqref{bss}.
Since $\| |x|^2 \nabla^\bot \psi \|_{L^\infty (\mathbb{R}^2)} \le C$, this term fulfills that
\[
\begin{split}
	&\int_0^t \int_{\mathbb{R}^2}
		\bigl| G_\alpha (t-s,x-y) \bigr|
		\bigl| |y|^{2+\alpha} y_j (\nabla \theta \cdot \nabla^\bot\psi) (s,y) \bigr|
	dyds\\
	&\le
	C\int_0^t
		\bigl\| G_\alpha (t-s) \bigr\|_{L^{p'} (\mathbb{R}^2)}
		\bigl\| |x|^\alpha x_j \nabla \theta \bigr\|_{L^p (\mathbb{R}^2)}
	ds.
\end{split}
\]
Here \eqref{drbdd} and \eqref{tokyu} yield that
\[
	\bigl\| |x|^\alpha x_j \nabla \theta \bigr\|_{L^p (\mathbb{R}^2)}
	\le
	\bigl\| |x|^{\alpha + 1} \nabla \theta \bigr\|_{L^p (\mathbb{R}^2)}
	\le
	\bigl\| |x|^2 \nabla \theta \bigr\|_{L^p (\mathbb{R}^2)}^{\frac{1+\alpha}2}
	\bigl\| \nabla \theta \bigr\|_{L^p (\mathbb{R}^2)}^{\frac{1-\alpha}2}
	\le
	C (1+t)^{- (\frac2\alpha (1-\frac1p)+\frac1\alpha)\frac{1-\alpha}2}.
\]
Concurrently, if we choose $p_1$ sufficiently near from $p$, then for $\frac1r = \frac1{p_1} - \frac1p + \frac12$,
\[
	\bigl\| (-\Delta)^{-1/2} (x\theta) \bigr\|_{L^{\frac{p_1 p}{p-p_1}} (\mathbb{R}^2)}
	\le
	C \bigl\| x\theta \bigr\|_{L^r (\mathbb{R}^2)}
	\le
	C \bigl\| |x|^2 \theta \bigr\|_{L^p (\mathbb{R}^2)}^{1/2} \bigl\| \theta \bigr\|_{L^{\frac{pr}{2p-r}} (\mathbb{R}^2)}^{1/2}
	\le C (1+t)^{\frac2\alpha (\frac1{p_1} - \frac1p)}.
\]
Therefore
\[
\begin{split}
	&\int_0^t \int_{\mathbb{R}^2}
		\bigl| G_\alpha (t-s) \bigr|
		\bigl| |x|^{2+\alpha} x_j \nabla \theta \cdot \nabla^\bot\psi (s) \bigr|
	dyds
	\le
	C \int_0^t
		(t-s)^{-\frac2{\alpha p}}
		(1+s)^{- (\frac2\alpha (1-\frac1p)+\frac1\alpha)\frac{1-\alpha}2}
	ds\\
	&+
	C \int_0^t
		(t-s)^{-\frac2{\alpha p_1}}
		(1+s)^{- (\frac2\alpha (1-\frac1p)+\frac1\alpha)\frac{1-\alpha}2+\frac2\alpha (\frac1{p_1}-\frac1p)}
	ds
	\le
	C (1+t).
\end{split}
\]
Similarly, for the second term on \eqref{bss}, we obtain that
\[
\begin{split}
	&\int_0^t \int_{\mathbb{R}^2}
		\bigl|  (x_j-y_j) \nabla G_\alpha (t-s,x-y) \bigr|
		\bigl| |y|^{2+\alpha} (\theta \nabla^\bot \psi)(s,y) \bigr|
	dyds\\
	&\le
	C \int_0^t
		\bigl\| x_j \nabla G_\alpha (t-s) \bigr\|_{L^{p'} (\mathbb{R}^2)} \bigl\| |x|^\alpha \theta \bigr\|_{L^p (\mathbb{R}^2)} 
	ds\\
	&\le
	C \int_0^t
		(t-s)^{-\frac2{\alpha p}} (1+s)^{-\frac2\alpha (1-\frac1p)+1}
	ds
	\le
	C (1+t).
\end{split}
\]
The fourth term on \eqref{bss} also is bouded by $C (1+t)$.
For the fifth term on \eqref{bss}, we see from \eqref{tokyu} that
\[
\begin{split}
	&\int_0^t
		\bigl\| |x|^{2+\alpha} G_\alpha (t-s) \bigr\|_{L^\infty (\mathbb{R}^2)}
		\bigl\| x_j (\nabla \theta \cdot \nabla^\bot\psi) (s) \bigr\|_{L^1 (\mathbb{R}^2)}
	ds\\
	&\le
	\int_0^t
		\bigl\| |x|^{2+\alpha} G_\alpha (t-s) \bigr\|_{L^\infty (\mathbb{R}^2)}
		\bigl\| x_j^2 \nabla \theta \bigr\|_{L^p (\mathbb{R}^2)}^{1/2}
		\bigl\| \nabla \theta \bigr\|_{L^p (\mathbb{R}^2)}^{1/2}
		\bigl\|  \nabla^\bot\psi \bigr\|_{L^{p'} (\mathbb{R}^2)}
	ds\\
	&\le
	C \int_0^t
		(t-s) (1+s)^{-\frac3{2\alpha}-\frac1{\alpha p}}
	ds
	\le
	C t.
\end{split}
\]
Applying those estimates to \eqref{bss}, we complete the proof.
\end{proof}

Finally, Lemma \ref{lem1} and Proposition \ref{prop-v2} show Theorem \ref{thm}.
Also we conclude Corollary \ref{cor} from \eqref{B-G}, Lemma \ref{lem2} and Proposition \ref{prop-v2}.

\end{document}